Grade 10Electric Current

A body of mass 2kg is released from a point 100m above the ground.calculate it's kinetic energy 80m on the point of release

Chiemerie
5 Years agoGrade 10

1 Answer

Harshit Singh
5 Years ago
Dear Student

V^2 = Vo^2 + 2g*(100-80) = 0 + 19.6*20 = 392,


KE = 0.5*M*V^2 = 0.5*2*392
= 392Joules.

Thanks